OpenStudy (anonymous):

@200205650 Choose the correct slope of the line that passes through the points (-4, 8) and (-3, -6). 7 -14 0 14

OpenStudy (anonymous):

Okay let me just do this on my calculator real quick.

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

Are these answer choices just 7x, -14x, 0x, and 14x?

OpenStudy (anonymous):

@jenny1003

OpenStudy (anonymous):

Yes i'M SORRY I wAS dOING sOMETHING

OpenStudy (anonymous):

It's okay. But are these answer choices just 7x, -14x, 0x, and 14x?

OpenStudy (anonymous):

@jenny1003

OpenStudy (sepeario):

Do you know how to find the slope of a line?

OpenStudy (sepeario):

The formula is \[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (sepeario):

So can you apply that equation to the two points and find the gradient?

OpenStudy (anonymous):

Oh ya now I remember how to do this. Ya sepeario reminded me of the formula. With that formula you can find the answer. :)

OpenStudy (anonymous):

\[\frac{ 8-(-6) }{ -4-(-3) }\]\[\frac{ 14 }{ -1 }\]\[-14\]

OpenStudy (anonymous):

Thank you very much @200205650